Let X be a discrete random variable. If Pr(X<9) = 2/9, and Pr(X<=9) = 6/18, then what is Pr(X=9)? Please specify your answer in decimal terms and round your answer to the nearest hundredth (e.g., enter 12 percent as 0.12
Answer #1

Solution:

Pr(X<9) = 2/9

Pr(X<=9) = 6/18

Given that  X is a discrete random variable.

Pr(X<=9) = Pr(X<9) + P(X = 9)

So ,

P(X = 9) = Pr(X<=9) -  Pr(X<9)  

= (6/18) - (2/9)

= 2/18

= 1/9

= 0.11

P(X = 9) = 0.11